The manger of a farmer's market has 500lb of gain that cost 1.20 per pound. How many pounds of meal costing .80 per pound should be mixed with the 500lb of grain to produce a mixture that cost 1.05 per pound?
Answer by mananth(16949)

--------price ---------------- quantity
Grain 1.2 ---------------- 500 lb
Meal 0.8 ---------------- x lb
Mixture 1.05 ---------------- 500 + x lb

1.2*500+0.8x=1.05(500+x)
600+0.8x=525+1.05x
0.8x-1.05x=-600 +525
-0.25 x = -75
/ -0.25
x = 300
300 lb of Meal
m.ananth@hotmail.ca